OpenStudy (anonymous):

Let F(x)=int_{0}^{x}(e)^{{(3t})^{4}} Find the Maclaurin polynomial polynomial of degree 5 for F(x)

OpenStudy (anonymous):

\[Let F(x)=\int\limits_{0}^{x}(e)^{{(3t})^{4}} \]

OpenStudy (anonymous):

that is what it looks like

OpenStudy (mathmate):

Integral of e^((3t)^4) dx. Start with Maclaurin's series for e^x: e^x = 1+x +x^2/2! + x^3/3! + x^4/4! + x^5/5! +.... for e^((3t)^4 substitute (3t)^4 for x: F(t) = 1 + (3t)^4 + ((3t)^4)^2/2! + ((3t)^4)^3/3! + ((3t)^4)^4/4! ... = 1 +81t^4 + (6561t^8)/2 +(177147t^12)/2 + (14348907t^16)/8 + (1162261467t^20)/40 + (31381059609t^24)/80 + ... You just need to integrate term by term as required.